OpenStudy (anonymous):

1. Find the distance between point A(-3, 5) and point B(4, -6) in the coordinate plane.

OpenStudy (anonymous):

\[d=\sqrt{(x-x1)^2 + (y-y1)^2}\]

OpenStudy (anonymous):

\[d = \sqrt{(-3+4)^{2} + (5-6)^{2}}\]

OpenStudy (anonymous):

D=\[\sqrt{(Y2-Y1)^2+(X2-X1)^2}\]

OpenStudy (kainui):

That's not just some random formula, it's the pythagorean theorem. Start thinking about points on a coordinate plane as really being triangles.

OpenStudy (anonymous):

when we have something to calculate why to waste time

OpenStudy (anonymous):

because if you think, you can just about do it in your head the distance between -3 and 4 is 7 the distance between -6 and 5 is 11 so you have a right triangle, one side is 7, the other sides is 11 and the hypotenuse (distance) is \[\sqrt{7^2+11^2}\]

OpenStudy (kainui):

Exactly. Plus, if you understand the concept, you can use that to help you learn new things or solve problems similar, which just knowing the formula can't help you with.

OpenStudy (anonymous):

good point